OpenStudy (anonymous):

Fans & Medals! Which functions have a positive rate of change and which have a negative rate of change? 1. 2x – 3y = 9 2. y = −4x + 5 3. x y −1 −5 1 1 3 7 4. x y 1 −2 2 −6 3 −10 A.positive B.negative

OpenStudy (anonymous):

@Nnesha

OpenStudy (anonymous):

@Destinymasha

OpenStudy (anonymous):

@Data_LG2

OpenStudy (anonymous):

rate is slope, slope is the value of m in y=mx+b \(\sf \Large m=\frac{rise}{run} =\frac{y_2-y_1}{x_2-x_1}\) For each choices, you have to determine the slope and check whether it is a negative slope or positive slope

OpenStudy (anonymous):

For 1, rearrange the equation into the form y=mx+b or evaluating the equation in terms of y

OpenStudy (anonymous):

so can you do this for now?

OpenStudy (anonymous):

try rearranging the equation into the for y=mx+b

OpenStudy (anonymous):

@LOLAri are you doing it?

OpenStudy (anonymous):

Yes I got positive for number 1 @Data_LG2

OpenStudy (anonymous):

so that is eliminated from your choices, how about the second one? it's already in the form y=mx+b

OpenStudy (anonymous):

One minute it took a while for me with the first one sooo lol